OpenStudy (anonymous):

The following system of equations is _____. y = -1/3 x + 2/3 2x + 6y = 4 consistent equivalent inconsistent

OpenStudy (mathmale):

You may find it easier to do this problem if you'd re-write 2x + 6Y = 4 in the form y=mx+b.

OpenStudy (vjt):

agreed^ once it is in that form it is much easier to work with

OpenStudy (mathmale):

Also, please be certain that you understand the meanings of "consistent," "inconsistent" and "equivalent."

OpenStudy (vjt):

agreed^ once it is in that form it is much easier to work with

OpenStudy (anonymous):

is it consisent?

OpenStudy (dayakar):

a1x+b1y+c1 =0 a2x+b2y+c2=0 are two equations write above equations in this form

OpenStudy (dayakar):